# Learning Unsigned Distance Fields from Local Shape Functions for 3D Surface Reconstruction

### v1

@@ 4.1 Experiment setup
Baselines. For our validation datasets, we compared our method against the state-of-the-art UDF learning models, which include unsupervised methods like CAP-UDF[17], LevelSetUDF[14], and DUDF[19], as well as the supervised learning method, GeoUDF[15]. We trained GeoUDF independently on different datasets to achieve optimal performance. Table. 1 shows the qualitative comparison between our methods and baselines. To evaluate performance, we calculate the Chamfer Distance (CD) and F1-Score (setting thresholds of 0.005 and 0.01) metrics between the ground truth meshes and the meshes extracted from the UDFs out by our model and each baseline model. For a fair comparison, we test all baseline models using the DCUDF[32] method. All experimental procedures are executed on NVIDIA RTX 4090 and A100 GPUs.
@@ 4.2 Experimental results
Synthetic data. For general 3D graphic models, ShapeNetCars, and DeepFashion3D, we obtain dense point clouds by randomly sampling on meshes.
@@
Considering that GeoUDF [15] is a supervised method, we retrain it on ShapeNetCars, and DeepFashion3D, which are randomly partitioned into training (70%), testing (20%), and validation subsets (10%). All models are evaluated in the validation sets, which remain unseen by any of the UDF learning models prior to evaluation. The first three rows of Fig. 5 show the visual comparison of reconstruction results, while Tab. 2 presents the quantitative comparison results of CD and F1-score. We test each method using their own mesh extraction technique, as shown in the inset figure, which display obvious visual artifacts such as small holes and non-smoothness. We thus apply DCUDF [32] , the state-of-art method, to each baseline model , extracting the surfaces as significantly higher quality meshes. Since our method utilizes DCUDF for surface extraction, we adopt it as the default technique to ensure consistency and fairness in comparisons with the baselines. Our method achieves stable results in reconstructing various types of surfaces, including both open and closed surfaces, and exhibits performance comparable to that of the SOTA methods. Noting that DUDF[19] requires normals during training, and GeoUDF utilizes the KNN approach to determine the nearest neighbors of the query points. Although DUDF and GeoUDF achieve better evaluations, they are less stable when dealing with point clouds with noise and outliers.
Clean Noise Outlier
CD $\downarrow$ F1 $\uparrow$ CD $\downarrow$ F1 $\uparrow$ CD $\downarrow$ F1 $\uparrow$
method $F1^{0.005}$ $F1^{0.01}$ $F1^{0.005}$ $F1^{0.01}$ $F1^{0.005}$ $F1^{0.01}$
ShapeNetCars [21] CAP-UDF [17] 2.432 0.523 0.888 2.602 0.194 0.381 4.982 0.183 0.314
LevelSetUDF [14] 1.534 0.561 0.908 2.490 0.209 0.401 4.177 0.199 0.363
GeoUDF [15] 1.257 0.571 0.889 1.232 0.351 0.873 4.870 0.187 0.346
DUDF [19] 0.568 0.903 0.991 3.180 0.312 0.527 4.235 0.168 0.308
Ours 1.085 0.510 0.938 1.114 0.427 0.922 1.272 0.485 0.771
DeepFashion3D [36] CAP-UDF [17] 1.660 0.417 0.818 1.892 0.336 0.542 4.941 0.172 0.430
LevelSetUDF [14] 1.500 0.403 0.856 1.488 0.453 0.729 4.328 0.203 0.468
GeoUDF [15] 0.652 0.864 0.977 1.258 0.380 0.957 4.463 0.147 0.300
DUDF [19] 0.381 0.991 0.998 1.894 0.334 0.535 4.970 0.144 0.272
Ours 0.932 0.652 0.983 1.150 0.361 0.976 1.029 0.549 0.973
Table 2: Quantitative evaluation of UDF learning methods (CD score is multiplied by 100).
@@
Noise & outliers. To evaluate our model with noisy inputs, we added Gaussian noise $\mathcal{N}(0,0.0025)$ to the clean data across all datasets for testing. The middle three rows in Fig. 5 display the reconstructed surface results from noisy point clouds, and Tab. 2 also presents the quantitative comparisons. It can be observed that our method can robustly reconstruct smooth surfaces from noisy point clouds. Additionally, we tested our method’s performance with outliers by converting 10% of the clean point cloud into outliers, as shown in the last three rows of Fig. 5. To further demonstrate the robustness of our method, we conducted experiments on point clouds with higher percentage of outliers. Our framework is able of reconstructing reasonable surfaces even with 50% outliers, as shown in Fig. 6. Furthermore, our approach can reconstruct high-quality geometry from point clouds containing both noise and outliers.
(a) Input (b) CAP-UDF (c) LevelSetUDF (d) GeoUDF (e) DUDF (f) Ours (g) GT
Figure 5: Visual comparisons on the synthetic dataset. First three rows: uniformly sampled points. Meddle three rows: point clouds with 0.25% added noise. Last three rows: point clouds with 10% outliers. All point clouds here have 48K points, except for the Bunny model, which has 100K points. We refer readers to the appendix for more visual results.
Figure 6: Our model demonstrates robustness to more outliers.
@@
Real-world scanned data. Dataset [38] provide several real-world scanned point clouds, as illustrated in Fig. 7 (Left), we evaluate our model on the dataset to demonstrate the effectiveness. Our approach can reconstruct smooth surfaces from scanned data containing noise and outliers. However, our model cannot address the issue of missing parts. This limitation is due to the local geometric training strategy, which is independent of the global shape. Additionally, we conduct tests on large scanned scenes to evaluate our algorithm, as shown in Fig. 7 (Right).
Figure 7: Reconstructed surfaces from real-world scanned point clouds.
4.3 Analysis & ablation studies
Efficiency. As a supervised UDF learning method, our approach has a significant improvement in training efficiency compared to GeoUDF [15].
Method Storage (GB) Data-prep (min) Training (h)
GeoUDF 120 0.5 36
Ours 0.59 0.02 14.5
As shown in the insert table, we calculate the data storage space required by GeoUDF when using ShapeNet as a training dataset. This includes the GT UDF values and point cloud data needed during the training process. Our synthetic point cloud patches training dataset occupies under 1GB, which is merely 0.5% of the storage needed for GeoUDF. Our network is very lightweight, with only 653KB of trainable parameters and a total parameter size of just 2MB. Additionally, we highlight time-saving benefits. The provided table illustrates the duration required to produce a single data sample for dataset preparation (“Data-prep”), as well as the total time for training (“Training”).
Patch radius. During the evaluation phase, the radius $r$ used to find the nearest points for each query point determines the size of the extracted patch and the range of effective query points in the space. As shown in Fig. 8, we analyzed the impact of different radii on the reconstruction results. An excessively small $r$ will generate artifacts, while an overly large $r$ will lose many details. In our experiments, we generally set $r$ to 0.018.
GT $r=0.08$ $r=0.10$ $r=0.20$ $r=0.30$
Figure 8: Comparison of different radii for extracting patches from the point cloud on reconstruction results.
@@ 4.3 Analysis & ablation studies
Denoising module. Our framework incorporates a denoising module to handle noisy point clouds. We conducted ablation experiments to verify the significance of this module. Specifically, we set $\lambda_{d}=0$ in the loss function Eq. (7) to disable the denoising module, and then retrained the network. As illustrated in Fig. 9, we present the reconstructed surfaces for the same set of noisy point clouds with and without the denosing module, respectively.
@@ 5 Conclusion
In this paper, we introduce a novel and efficient neural framework for surface reconstruction from 3D point clouds by learning UDFs from local shape functions. Our key insight is that 3D shapes exhibit simple patterns within localized regions, which can be exploited to create a training dataset of point cloud patches represented by mathematical functions. As a result, our method enables efficient and robust surfaces reconstructions without the need for shape-specific training. Extensive experiments on various datasets have demonstrated the efficacy of our method. Moreover, our framework achieves superior performance on point clouds with noise and outliers.
Limitations & future work. Owing to its dependence solely on local geometric features, our approach fails to address tasks involving incomplete point cloud reconstructions. However, as a lightweight framework, our model can readily be integrated into other unsupervised methods to combine the global features with our learned local priors. Furthermore, in our future work, we intend to design a method that dynamically adjusts the radius based on local feature sizes [39] of 3D shapes when extracting local point cloud patches for queries, aiming to improve the accuracy of the reconstruction.
Appendix A Appendix
A.1 More results
As shown in Fig. 10 and Fig. 11, we provide more visual comparisons on the DeepFashion3D and ShapeNetCars dataset, using point clouds containing noise and outliers.
(a) Input (b) CAP-UDF (c) LevelSetUDF (d) GeoUDF (e) DUDF (f) Ours (g) GT
Figure 10: More visual results on the DeepFashion3D dataset. Top three rows: Reconstruction results under noise-free conditions. Bottom three rows: Reconstruction results under noise condition.
(a) Input (b) CAP-UDF (c) LevelSetUDF (d) GeoUDF (e) DUDF (f) Ours (g) GT
Figure 11: More visual results on the synthetic datasets with outliers.

### v2

@@ 4.1 Setup
Baselines & metrics. For our validation datasets, we compared our method against the state-of-the-art UDF learning models, which include unsupervised methods like CAP-UDF [48], LevelSetUDF [49], and DUDF [15], as well as the supervised learning method, GeoUDF [37]. We trained GeoUDF independently on different datasets to achieve optimal performance. Table 1 shows the qualitative comparison between our methods and baselines. To evaluate performance, we compare our approach with other baseline models in terms of $L_{1}$-Chamfer Distance (CD), F1-Score (setting thresholds of 0.005 and 0.01), and normal consistence (NC) metrics between the ground truth meshes and the meshes extracted from learned UDFs. For a fair comparison, we adopt the same DCUDF [18] method for mesh extraction. All experiments are conducted on NVIDIA RTX 4090 GPU.
@@
Input(48K)
GT
CAP-UDF
LevelSetUDF
GeoUDF
DUDF
Ours
     Clean Noise(0.25%) Outliers(10%)
Figure 5: Visual comparisons of reconstruction results on the synthetic dataset. We provide more results in the supplementary materials.
Figure 6: We compare other methods using their own mesh extraction techniques against the DCUDF approach.
@@ 4.2 Experimental results
Synthetic data. For general 3D graphic models, ShapeNetCars, and DeepFashion3D, we obtain dense point clouds by randomly sampling on meshes. Considering that GeoUDF [37] is a supervised method, we retrain it on ShapeNetCars, and DeepFashion3D, which are randomly partitioned into training (70%), testing (20%), and validation subsets (10%). All models are evaluated in the validation sets, which remain unseen by any of the UDF learning models prior to evaluation. Figure 5 illustrates the visual comparison of reconstruction results, and Table 3 presents the quantitative comparison in terms of evaluation metrics. We test each method using their own mesh extraction technique, as shown in Fig. 6, which display obvious visual artifacts such as small holes and non-smoothness. We thus apply DCUDF [18] , the state-of-art method, to each baseline model , extracting the surfaces as significantly higher quality meshes. Since our method utilizes DCUDF for surface extraction, we adopt it as the default technique to ensure consistency and fairness in comparisons with the baselines. Our method achieves stable results in reconstructing various types of surfaces, including both open and closed surfaces, and exhibits performance comparable to that of the SOTA methods. Noting that DUDF [15] requires normals during training, and GeoUDF utilizes the KNN approach to determine the nearest neighbors of the query points. As a result, DUDF and GeoUDF are less stable when dealing with point clouds with noise and outliers, as shown in Fig. 5.
Table 2: We compare our method with other UDF learning methods in terms of $L_{1}$-Chamfer distance (×$\times 100$), F-score with thresholds of 0.005 and 0.01, and normal consistence. The best results are highlighted with 1st and 2nd.
Clean Noise Outliers
Method CD $\downarrow$ $F1^{0.005}$ $F1^{0.01}$ NC $\uparrow$ CD $\downarrow$ $F1^{0.005}$ $F1^{0.01}$ NC $\uparrow$ CD $\downarrow$ $F1^{0.005}$ $F1^{0.01}$ NC $\uparrow$
ShapeNetCars [8] CAP-UDF [48] 1.324 0.451 0.884 0.945 1.208 0.286 0.632 0.833 2.918 0.194 0.304 0.725
LevelSetUDF [49] 1.223 0.464 0.892 0.950 1.176 0.299 0.654 0.842 2.763 0.200 0.379 0.753
GeoUDF [37] 0.832 0.562 0.901 0.947 1.869 0.205 0.608 0.847 2.598 0.144 0.327 0.757
DUDF [15] 0.773 0.697 0.871 0.945 1.939 0.214 0.508 0.826 2.949 0.135 0.329 0.722
Ours 0.611 0.645 0.927 0.967 0.614 0.546 0.913 0.966 0.705 0.437 0.751 0.944
DeepFashion3D [24] CAP-UDF [48] 0.885 0.353 0.809 0.973 0.803 0.375 0.719 0.871 5.150 0.173 0.425 0.863
LevelSetUDF [49] 0.881 0.335 0.848 0.983 0.723 0.384 0.750 0.874 5.083 0.278 0.430 0.882
GeoUDF [37] 0.625 0.525 0.957 0.973 0.676 0.506 0.951 0.882 3.883 0.146 0.310 0.764
DUDF [15] 0.506 0.586 0.972 0.965 0.934 0.333 0.530 0.849 2.558 0.118 0.265 0.821
Ours 0.548 0.565 0.966 0.988 0.614 0.541 0.967 0.906 0.559 0.463 0.966 0.909
Table 3: Comparison of time efficiency. We measured the average runtime in minutes. "#Params" denotes the number of network parameters, while "Size" refers to the storage space occupied by these parameters.
Method SRB DeepFashion3D ShapeNetCars #Param Size (KB)
CAP-UDF 15.87 10.5 10.6 463100 1809
LoSF + CAP-UDF 6.84 4.32 4.40 - -
LevelSetUDF 15.08 13.65 14.67 463100 1809
LoSF + LevelSetUDF 6.13 4.85 4.97 - -
DUDF 14.28 11.12 14.58 461825 1804
GeoUDF 0.08 0.07 0.07 253378 990
Ours 0.87 0.51 0.42 167127 653
@@
Noise & outliers. To evaluate our model with noisy inputs, we added Gaussian noise $\mathcal{N}(0,0.25\%)$ to the clean data across all datasets for testing. The middle three columns in Fig. 5 display the reconstructed surface results from noisy point clouds, and Table 3 also presents the quantitative comparisons. It can be observed that our method can robustly reconstruct smooth surfaces from noisy point clouds. Additionally, we tested our method’s performance with outliers by converting 10% of the clean point cloud into outliers, as shown in the last three columns of Fig. 5. Experimental results demonstrate that our method can handle up to 50% outliers while still achieving reasonable results. Even in the presence of both noise and outliers, our method maintains a high level of robustness. The corresponding results are provided in the supplementary materials.
@@
Real scanned data. Dataset [3] provide several real scanned point clouds, as illustrated in Fig. 7, we evaluate our model on the dataset to demonstrate the effectiveness. Our approach can reconstruct smooth surfaces from scanned data containing noise and outliers. However, our model cannot address the issue of missing parts. This limitation is due to the local geometric training strategy, which is independent of the global shape.
Input CAP-UDF LevelSetUDF DUDF GeoUDF Ours Method CAP-UDF LevelSetUDF DUDF GeoUDF Ours CD($\times$100) $\downarrow$ 2.101 1.637 4.207 1.762 1.551 $F1^{0.005}$uparrow 0.410 0.464 0.378 0.557 0.589 $F1^{0.01}$uparrow 0.657 0.730 0.699 0.849 0.878 NC $\uparrow$ 0.934 0.943 0.907 0.962 0.963
Figure 7: Evaluations on real scanned data. The evaluation metrics presented in the table represent the average results of these models.
4.3 Analysis
Efficiency. We compare the time complexity of our method with other methods, as shown in Tab. 3. All tests were conducted on an Intel i9-13900K CPU and an NVIDIA RTX 4090 GPU. Computational results show that supervised, local feature-based methods like our approach and GeoUDF [37] significantly outperform unsupervised methods in terms of computational efficiency. Additionally, our method has a significant improvement in training efficiency compared to GeoUDF. Utilizing ShapeNet as the training dataset, GeoUDF requires 120GB of storage space. In contrast, our method employs a shape-category-independent dataset, occupying merely 0.50GB of storage. Our network is very lightweight, with only 653KB of trainable parameters and a total parameter size of just 2MB. Compared to GeoUDF, which requires 36 hours for training, our method only requires 14.5 hours.
Patch radius and point density. During the evaluation phase, the radius $r$ used to find the nearest points for each query point determines the size of the extracted patch and the range of effective query points in the space. The choice of radius directly influences the complexity of the geometric features captured. When normalizing point clouds to a unit bounding box, we set the radius, $r=0.018$. This setting achieves satisfactory reconstruction for our testing datasets. In the supplementary materials, we present a bias analysis experiment comparing the synthesized local patches and the local geometries extracted from the test point cloud data. The experimental results confirm that setting $r$ to 0.018 maintains a relatively low bias, suggesting its effectiveness. Users can conduct a preliminary bias analysis based on our well-trained model to adjust the size of the radius according to the complexity of the input point cloud. This process is not time-consuming. Through experimental testing (refer to the supplementary materials), our algorithm ensures reasonable reconstruction, provided that there are at least 30 points within a unit area. A possible way for mitigating issues arising from low sampling rates is to apply an upsampling module [37] during the pre-processing step.
Input Ours Ablation (a) Ablation (b)
Figure 8: Ablation studies on Cross-Attn module. (a) Without the Points-Net and Cross-Attn modules. (b) Without the Cross-Attn module. (CD score is multiplied by 100)
4.4 Ablation studies
Cross-Attn module. Our main goal is to derive the UDF value for a query point by learning the local geometry within a radius $r$. To achieve this, we utilize Points-Net to capture the point cloud features $\mathbf{f}_{p}$ of local patches. This process enables the local geometry extracted from test data to align with the synthetic data through feature matching, even in the presence of noise or outliers. Vectors-Net is tasked with learning the features $\mathbf{f}_{v}$ of the set of vectors pointing towards the query point, which includes not only the position of the query point but also its distance information. The Cross-Attn module then processes these local patch features $\mathbf{f}_{p}$ as keys and values to query the vector features $\mathbf{f}_{v}$, which contain distance information, returning the most relevant feature $\mathbf{f}_{G}$ that determines the UDF value. See Fig. 8 for two ablation studies on noisy point clouds.
@@ 4.4 Ablation studies
Denoising module. Our framework incorporates a denoising module to handle noisy point clouds. We conducted ablation experiments to verify the significance of this module. Specifically, we set $\lambda_{d}=0$ in the loss function to disable the denoising module, and then retrained the network. As illustrated in Fig. 9, we present the reconstructed surfaces for the same set of noisy point clouds with and without the denosing module, respectively.
@@
4.5 Results of unsupervised integration
Our LoSF-UDF approach offers better initialization for unsupervised methods, including CAP-UDF [48], LevelSetUDF [49], and DEUDF [13]. We evaluated 12 models on the Threescan dataset [21], each containing rich details. Using the integration framework based on our lightweight LoSF-UDF, we achieve comparable or even superior reconstruction results, as illustrated in Fig. 10 and Tab. 4. More importantly, we improve the efficiency of original unsupervised methods as shown in Tab. 3. Considering the DEUDF is not open-source, we employ their proposed loss functions to train a SIREN network [39] on our own. For the loss function terms that require normal information, we used the method of PCA [17] to estimate the normals during the optimization process. Thanks to the robustness of LoSF, the accuracy of its estimation has been enhanced.
Table 4: Quantitative comparison results of the integrated framework with unsupervised methods.
Method LoSF *+CAP-UDF [48] *+LevelSetUDF [49] *+SIREN [45]
CD($\times$100) $\downarrow$ 0.409 0.447 0.429 0.217
$F1^{0.005}$uparrow 0.638 0.609 0.610 0.906
$F1^{0.01}$uparrow 0.958 0.936 0.952 0.984
NC $\uparrow$ 0.964 0.946 0.962 0.969
Input LoSF *+CAP-UDF *+LevelSetUDF *+SIREN
Figure 10: Reconstruction results of the integrated framework.
@@ 5 Conclusion
In this paper, we introduce a novel and lightweight neural framework for surface reconstruction from 3D point clouds by learning UDFs from local shape functions. Our key insight is that 3D shapes exhibit simple patterns within localized regions, which can be exploited to create a training dataset of point cloud patches represented by mathematical functions. As a result, our method enables efficient and robust surfaces reconstruction without the need for shape-specific training, even in the presence of noise and outliers. Extensive experiments on various datasets have demonstrated the efficacy of our method. Moreover, our lightweight framework can be integrated with unsupervised methods to provide rapid and reliable initialization, enhancing both efficiency and accuracy.
Acknowledgement
The NTU authors were supported in part by the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22) and the RIE2020 Industry Alignment Fund–Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). The DLUT authors were supported by the National Natural Science Foundation of China under Grants 62402083 and T2225012, the National Key R&D Program of China under Grants 2021YFA1003003. F. Hou was supported by the Basic Research Project of ISCAS (ISCAS-JCMS-202303) and the Major Research Project of ISCAS (ISCAS-ZD-202401). J. Hou was supported by the Hong Kong RGC under Grants 11219422 and 11219324.